# Mind the Gap: From Resolving Theoretical Foundations of Chiral(ity)-Induced Spin Selectivity to Pioneering Implementations in Quantum Sensing

### v2

@@ II.3 The Role of Symmetries: Chiral-Induced or Chirality-Induced?
In light of both geometric- and point-chirality contributing to CISS despite the different analyses of their SOC-contributions (the former accountable purely from effective helical-motion, while the latter is better described as emergent from the configuration of $\sigma$- and $\pi$-type orbital overlaps), it is natural to probe at a higher level and question if chirality by itself can give rise to any relevant topological electronic properties [99] in CISS. At its simplest, we can map a 1D helix to a toy Su-Schrieffer-Heeger-like model by assuming continuous screw-sense symmetry and coherent transport. When the sites are arranged such that they admit a perfectly periodic unit-cell, the conservation of screw-sense helical momentum labels the model with a Zak phase upon the introduction of an external magnetic field (that breaks time-reversal symmetry), giving rise to a quantised charge-pump that carries no net spin; see Sacramento & Madeira’s [100] study for a deeper investigation into the topological properties of such 1D chiral electronic chains. Any spin-selective effect that does not invoke an explicit magnetic field should therefore arise beyond the single Bloch-band description, for instance by employing multi-orbital frameworks to remain within the purely coherent paradigm as demonstrated by Utsumi et al. [73, 72]. Under the premise that a chiral biomolecule primarily functions as an orbital-momentum selector due to its weak SOC, Liu et al. go beyond a molecule-only picture and reveal the presence of a topological orbital texture via which orbital-polarisation is generated and then converted to spin-polarisation at the interface with a substrate of non-negligible SOC [70]. However, the applicability of either coherent-transport approaches to explain experimental findings is uncertain: initial calculations by Gersten et al. [101] and recently experimental investigations from Li et al. [102] using both nonmagnetic and magnetic break-junction setups found that no significant dependence of magnetoconductance or I-V characteristics on enantiomeric handedness within the fully coherent transport regime, owing to interfacial SOC being too weak and transit time being too short for spin-filtering to occur at a coherent time-scale, accentuating the setup-dependent nature for the origin of the CISS effect.
More generally, in the absence of external magnetic fields and any crystalline structure, Weyl fermions with finite Chern number appear at time-reversal-invariant momenta for all symmorphic chiral space groups [103]. Unlike Weyl fermions generated by band inversion, these chiral fermions exist a priori and their finite chiral charge implies that they can only be destroyed via pairwise annihilation, requiring application of a large magnetic field. Whether they manifest in chiral soft-matter is an intriguing avenue for further work, but it suffices to note at current that chirality is not a sufficient factor for spin-selectivity.
@@ III.2 CISS in Excited-States: Scattering Models for Photoexcitation
Given the complications as presented above, one prominent theoretical goal has been to find some analytical description that can capture the essential physics of CISS-transmission. A foundational effort is given by Yeganeh et al. [114], in which scattering theory is used to model electron transmission through a helical potential dressed with a spin-orbit coupling field, assuming elastic first-order scattering. Their results yield that a finite pitch result in differential angular momentum transfer, validating chirality’s role in spin-selectivity. However, asymmetry readings are too low to match experimental results, rendering chirality alone to be a necessary but insufficient factor. Subsequent work expands on this approach by considering the multiple-scattering regime [115], interfacial conversion of orbital-to-spin angular momentum due to substrate SOC [101], inelastic effects [116], and a minimal model that generalizes the CISS-helix into a CISS-dipole [79]. However, these analytic methods either do not give rise to anomalous spin-polarization readings as measured in Mott polarimetry, or cannot (in the case of interfacial effects) account for similar spin-polarization readings when substrates with low-SOC are used [117]. It stands to reason that either some non-equilibrium [118]mechanism perturbs the set-up significantly that is not captured by the above methods, or there is some omitted factor in the chiral molecule itself that is responsible. Could there be more to the helical contribution to CISS than currently treated? We now move to discussing CISS-transport in waveguides in the next subsection, where we cover efforts to find some rigorous starting-point to study CISS.
@@ III.3 CISS in Waveguides: Continuum Approximations and Geometric Efforts
where the electric field is augmented as $\vec{E}^{\prime}=\vec{E}+(\vec{\Omega}\times\vec{r})\times\vec{B}$, accounting for a Lorentz boost due to an external magnetic field. Since SOC is to order $m^{-2}$ while the spin-rotating term is to order $m^{-1}$, SOC contributions can be considered small in vacuum conditions; and in the absence of an external magnetic field, the spin-rotation coupling, $\Omega\cdot\sigma$, provides spin-splitting. Though this analysis corroborates the existence of geometric effects in CISS, their strength remains a matter of contention. On one hand, heuristically accounting for a linear electronic momentum of 1eV as with Gersten et al. [101] makes such effects negligible. On the other hand, recent work from Bradbury et al. [11] account for non-Born-Oppenheimer effects using a phase-space formulation of electronic structure, predicting a strong spin-Coriolis effect even for electrons of small molecules. And while there are innovative schemes to explicitly test curvature-dependence of CISS in experiments[126], much of biomolecular-CISS as observed is a non-equilibrium phenomenon where the effects of electron correlations and electron structure in materials cannot be neglected. Accounting for these likely-more-significant contributions calls for a different approach.
@@ III.4 CISS in Waveguides: Coarse-Grained Models
Kane-Mele-like $\sum_{m}i\lambda\psi_{m}^{\dagger}(\vec{\delta}_{1}\times\vec{\delta}_{2})\cdot\mathbf{\sigma}\psi_{m+2}+\mathrm{H.c.}$
@@
$t\sum_{m}\psi_{m}^{\dagger}e^{i\lambda(d_{m,m+1}\times\mathbf{E}_{m,m+1})\cdot\sigma}\psi_{m+1}+\mathrm{H.c.}$
Table 1: Phenomenological SOC expressions
@@ III.7 CISS in Spin Chemistry and Electron Transfer Reactions
In addition to spin filtering mechanisms, spin coherent mechanisms have also been proposed. Luo & Hore[179] formulated a phenomenological model, to investigate CISS effects in cryptochrome magnetoreception, that assumes spin polarization during the formation and recombination of a singlet born radical pair. Formally they arrive at the initial state
@@
Another spin coherent mechanism was introduced by Fay [184] who demonstrated that, based on a microscopic derivation for a one-step electron transfer in radical pairs, chirality in conjunction with SOC generates quantum coherence between singlet and triplet spin-states, yet does not generate spin polarization; we briefly detail the model as illustrated in Fig. 18.
Consider a charge-transfer reaction for a photoexcited radical pair system. The excited precursor state can be designated $\mathinner{|{0}\rangle}=\mathinner{|{DA^{*}}\rangle}$, and the charge-separated radical-pair state can be designated $\mathinner{|{1}\rangle}=\mathinner{|{D^{\bullet+}A^{\bullet-}}\rangle}$, where DD(AA) indicates the donor(acceptor). Of note, $\mathinner{|{0}\rangle}$ only exists in the singlet configuration while $\mathinner{|{1}\rangle}$ can exist in either singlet and triplet configuration. The Hamiltonian for such a system dressed in an SOC field is thus:
@@
Figure 20: The effect of CISS on the spin states of spin-correlated radical pairs (SCRPs), as revealed by time-resolved electron paramagnetic resonance (TREPR). (A) In the absence of CISS, a SCRP formed in a singlet state evolves in the presence of a strong static magnetic field $\bf{B}_{0}$. (B, E) Alignment of the chiral molecules parallel ($\theta=0^{\circ},180^{\circ}$) or perpendicular ($\theta=90^{\circ}$) to $\bf{B}_{0}$. (C, F) TREPR spectra showing singlet-born SCRP signals with and without CISS contributions, where red traces indicate full CISS generated spin polarization, black traces indicate partial CISS contributions, and blue the case without CISS. (D, G) Schematic diagrams of spin states under 100% CISS contribution. $D$ represents spin-spin interactions. Reprinted with permission from Eckvahl et al., Science 382, 197–201 (2023). [39]. Copyright 2023 Authors.
@@ IV.1 CISS-Inspired Sensing and Detection
In fact, from a quantum thermodynamic perspective, this implies that the state transformation mechanisms underlying the CISS effect could be acting as passive isothermal quantum erasers or initializers that are robust at room-temperature. In treating CISS as a quantum channel encoding spin-selective transmission amplitudes, chiral asymmetry due to molecular geometry, and open quantum systems effects like dephasing [157, 131], one can conceptualise formally recasting it as an energetically optimal qubit transformer [122, 218], thereby putting it in the same family as Landauer erasers, Maxwell’s demon type quantum engines, and information-powered refrigerators.
@@ IV.2 Spin-Correlated Radical Pair Platforms
Recently, Lin et al. [22] introduced an emerging platform based on donor-chiral bridge-acceptor (D-$\chi$-A) molecules that boast significant magnetic field effects (MFEs) on fluorescence intensity and SCRP lifetimes, even in solution at ambient conditions. These systems build upon previous designs by the Mani group for donor-bridge-acceptor (D-B-A) molecules for optical readout of magnetic-field-sensitive SCRPs at room temperature. While the reported changes for those systems were modest, the D-$\chi$-A design exhibit significant magnetic field effects on recombination fluorescence, where the chiral bridge provides sufficient electronic couplings to ensure long-lived SCRPs. Although the role of CISS was not explicitly resolved by Lin et al., similar chiral-bridged molecules have been used to investigate and demonstrate the CISS effect, giving credence to the plausibility that CISS could contribute in these systems [260, 39]. The advantage of molecular synthetic control, was demonstrated through three structural modifications to the donor site via distance extension, torsional locking, and planarization. These changes produced an observed magnetic sensitivity increase of $200\%$ in emission lifetimes, with up to $30\%$ modulation of total emission intensity (Fig. 24). Furthermore, as these structural changes allow the system rigidity and donor-acceptor distance to be altered, and spin relaxation via singlet-triplet dephasing mechanisms to be modulated, it is plausible that some of the enhancement and control schemes identified in biological SCRPs could be implemented [261, 262, 249].
D-$\chi$-A SCRPs represent a promising emerging platform of SCRP-based molecular quantum sensors [263], with the potential to extend beyond classical sensing capabilities to address a wide range of applications from biological to condensed matter systems [264, 265, 266]. Their long-lived spin coherence, room-temperature operation, optical addressability, and structural tunability positions them as strong candidates for developing type-II and type-III quantum sensors that exploit quantum resources such as quantum coherence and entanglement. Moreover, these systems offer a complementary perspective by not only acting as sensors, but also providing testbeds, alongside NV-based quantum sensing to probe fundamental questions about quantum properties of radical pairs and the role of CISS [267, 53, 268].

### v3

@@ II.3 The Role of Symmetries: Chiral-Induced or Chirality-Induced?
In light of both geometric- and point-chirality contributing to CISS despite the different analyses of their SOC-contributions (the former accountable purely from effective helical-motion, while the latter is better described as emergent from the configuration of $\sigma$- and $\pi$-type orbital overlaps), it is natural to probe at a higher level and question if chirality by itself can give rise to any relevant topological electronic properties [102] in CISS. At its simplest, we can map a 1D helix to a toy Su-Schrieffer-Heeger-like model by assuming continuous screw-sense symmetry and coherent transport. When the sites are arranged such that they admit a perfectly periodic unit-cell, the conservation of screw-sense helical momentum labels the model with a Zak phase that is only quantized when time-reversal symmetry is broken (e.g. by an external magnetic field); see Sacramento & Madeira’s [103] study for a deeper investigation into the topological properties of such 1D chiral electronic chains. Any spin-selective effect that does not invoke an explicit magnetic field should therefore arise beyond the single Bloch-band description, for instance by employing multi-orbital frameworks to remain within the purely coherent paradigm as demonstrated by Utsumi et al. [76, 75]. Under the premise that a chiral biomolecule primarily functions as an orbital-momentum selector due to its weak SOC, Liu et al. go beyond a molecule-only picture and reveal the presence of a topological orbital texture via which orbital-polarisation is generated and then converted to spin-polarisation at the interface with a substrate of non-negligible SOC [73]. However, the applicability of either coherent-transport approaches to explain experimental findings is uncertain: initial calculations by Gersten et al. [104] and recently experimental investigations from Li et al. [105] using both nonmagnetic and magnetic break-junction setups found that no significant dependence of magnetoconductance or I-V characteristics on enantiomeric handedness within the fully coherent transport regime, owing to interfacial SOC being too weak and transit time being too short for spin-filtering to occur at a coherent time-scale, accentuating the setup-dependent nature for the origin of the CISS effect. More generally, in the absence of external magnetic fields and any crystalline structure, Weyl fermions with finite Chern number appear at time-reversal-invariant momenta for all symmorphic chiral space groups [106]. Unlike Weyl fermions generated by band inversion, these chiral fermions exist a priori and their finite chiral charge implies that they can only be destroyed via pairwise annihilation, requiring application of a large magnetic field. Whether they manifest in chiral soft-matter is an intriguing avenue for further work, but it suffices to note at current that chirality is not a sufficient factor for spin-selectivity.
@@ III.2 CISS in Excited-States: Scattering Models for Photoexcitation
Given the complications presented above, a prominent goal has been to find some analytical description that captures the essential physics of CISS-transmission. A foundational effort is given by Yeganeh et al. [117], in which scattering theory is used to model electron transmission through a helical potential dressed with a spin-orbit coupling field, assuming elastic first-order scattering. Their results yield that a finite pitch result in differential angular momentum transfer, validating chirality’s role in spin-selectivity. However, asymmetry readings are too low to match experimental results, rendering chirality alone to be a necessary but insufficient factor. Subsequent work expands on this approach by considering the multiple-scattering regime [118], interfacial conversion of orbital-to-spin angular momentum due to substrate SOC [104], inelastic effects [119], and a minimal model that generalizes the CISS-helix into a CISS-dipole [82]. However, these analytic methods either do not give rise to anomalous spin-polarization readings as measured in Mott polarimetry, or cannot (in the case of interfacial effects) account for similar spin-polarization readings when substrates with low-SOC are used [120]. It stands to reason that either some non-equilibrium [121]mechanism perturbs the set-up significantly that is not captured by the above methods, or there is some omitted factor in the chiral molecule itself that is responsible. Could there be more to the helical contribution to CISS than currently treated? We now move to discussing CISS-transport in waveguides in the next subsection, where we cover efforts to find some rigorous starting-point to study CISS.
@@ III.3 CISS in Waveguides: Continuum Approximations and Geometric Efforts
where the electric field is augmented as $\vec{E}^{\prime}=\vec{E}+(\vec{\Omega}\times\vec{r})\times\vec{B}$, accounting for a Lorentz boost due to an external magnetic field. Note that the $\lambda$ denotes $\hbar^{2}/(4m_{0}^{2}c^{2})$ but is used more generally in the literature to refer to phenomenological tight-binding SOC strength. Since SOC is to order $m^{-2}$ while the spin-rotating term is to order $m^{-1}$, SOC contributions can be considered small in vacuum conditions; and in the absence of an external magnetic field, the spin-rotation coupling, $\Omega\cdot\sigma$, provides spin-splitting. Though this analysis corroborates the existence of geometric effects in CISS, their strength remains a matter of contention. On one hand, heuristically accounting for a linear electronic momentum of 1eV as with Gersten et al. [104] makes such effects negligible. On the other hand, recent work from Bradbury et al. [11] account for non-Born-Oppenheimer effects using a phase-space formulation of electronic structure, predicting a strong spin-Coriolis effect even for electrons of small molecules. And while there are innovative schemes to explicitly test curvature-dependence of CISS in experiments[129], much of biomolecular-CISS as observed is a non-equilibrium phenomenon where the effects of electron correlations and electron structure in materials cannot be neglected. Accounting for these likely-more-significant contributions calls for a different approach.
@@ III.4 CISS in Waveguides: Coarse-Grained Models
Kane-Mele-like $\sum_{m}i\lambda\psi_{m}^{\dagger}(\vec{d}_{m,m+1}\times\vec{d}_{m+1,m+2})\cdot\mathbf{\sigma}\psi_{m+2}+\mathrm{H.c.}$
@@
$t\sum_{m}\psi_{m}^{\dagger}e^{i\lambda(\vec{d}_{m,m+1}\times\mathbf{E}_{m,m+1})\cdot\sigma}\psi_{m+1}+\mathrm{H.c.}$
Table 1: Phenomenological SOC expressions. $\lambda$ denotes the phenomenological SOC strength; $\vec{d}_{m,n}$ denotes the bond vector from site m to n, while $E_{m,n}$ denotes the local electric field on that bond.
@@ III.7 CISS in Spin Chemistry and Electron Transfer Reactions
Another spin coherent mechanism was introduced by Fay [187] who demonstrated that, based on a microscopic derivation for a one-step electron transfer in radical pairs, chirality in conjunction with SOC generates quantum coherence between singlet and triplet spin-states, yet does not generate spin polarization; we briefly detail the model as illustrated in Fig. 18.Consider a charge-transfer reaction for a photoexcited radical pair system. The excited precursor state can be designated $\mathinner{|{0}\rangle}=\mathinner{|{DA^{*}}\rangle}$, and the charge-separated radical-pair state can be designated $\mathinner{|{1}\rangle}=\mathinner{|{D^{\bullet+}A^{\bullet-}}\rangle}$, where DD(AA) indicates the donor(acceptor). Of note, $\mathinner{|{0}\rangle}$ only exists in the singlet configuration while $\mathinner{|{1}\rangle}$ can exist in either singlet and triplet configuration. The Hamiltonian for such a system dressed in an SOC field is thus:
@@
Figure 20: The effect of CISS on the spin states of spin-correlated radical pairs (SCRPs), as revealed by time-resolved electron paramagnetic resonance (TREPR). (A) In the absence of CISS, a SCRP formed in a singlet state evolves in the presence of a strong static magnetic field $\bf{B}_{0}$. (B, E) Alignment of the chiral molecules parallel ($\theta=0^{\circ},180^{\circ}$) or perpendicular ($\theta=90^{\circ}$) to $\bf{B}_{0}$. (C, F) TREPR spectra showing singlet-born SCRP signals with and without CISS contributions, where red traces indicate full CISS generated spin polarization, black traces indicate partial CISS contributions, and blue the case without CISS. (D, G) Schematic diagrams of spin states under 100% CISS contribution. $D$ represents spin-spin interactions. Reprinted with permission from Eckvahl et al.[40], Science 382, 197–201 (2023). Copyright 2023 Authors.
@@ IV.1 CISS-Inspired Sensing and Detection
In fact, from a quantum thermodynamic perspective, this implies that the state transformation mechanisms underlying the CISS effect could be acting as passive isothermal quantum erasers or initializers that are robust at room-temperature. In treating CISS as a quantum channel encoding spin-selective transmission amplitudes, chiral asymmetry due to molecular geometry, and open quantum systems effects like dephasing [160, 134], one can conceptualise formally recasting it as an energetically optimal qubit transformer [125, 221], thereby putting it in the same family as Landauer erasers, Maxwell’s demon type quantum engines [222], and information-powered refrigerators.
@@ IV.2 Spin-Correlated Radical Pair Platforms
Recently, Lin et al. [23] introduced an emerging platform based on donor-chiral bridge-acceptor (D-$\chi$-A) molecules that boast significant magnetic field effects (MFEs) on fluorescence intensity and SCRP lifetimes, even in solution at ambient conditions. These systems build upon previous designs by the Mani group for donor-bridge-acceptor (D-B-A) molecules for optical readout of magnetic-field-sensitive SCRPs at room temperature. While the reported changes for those systems were modest, the D-$\chi$-A design exhibit significant magnetic field effects on recombination fluorescence, where the chiral bridge provides sufficient electronic couplings to ensure long-lived SCRPs. Although the role of CISS was not explicitly resolved by Lin et al., similar chiral-bridged molecules have been used to investigate and demonstrate the CISS effect, giving credence to the plausibility that CISS could contribute in these systems [264, 40]. The advantage of molecular synthetic control, was demonstrated through three structural modifications to the donor site via distance extension, torsional locking, and planarization. These changes produced an observed magnetic sensitivity increase of $200\%$ in emission lifetimes, with up to $30\%$ modulation of total emission intensity (Fig. 24). Furthermore, as these structural changes allow the system rigidity and donor-acceptor distance to be altered, and spin relaxation via singlet-triplet dephasing mechanisms to be modulated, it is plausible that some of the enhancement and control schemes identified in biological SCRPs could be implemented [265, 266, 253]. D-$\chi$-A SCRPs represent a promising emerging platform of SCRP-based molecular quantum sensors [267], with the potential to extend beyond classical sensing capabilities to address a wide range of applications from biological to condensed matter systems [268, 269, 270]. Their long-lived spin coherence, room-temperature operation, optical addressability, and structural tunability positions them as strong candidates for developing type-II and type-III quantum sensors that exploit quantum resources such as quantum coherence and entanglement. Moreover, these systems offer a complementary perspective by not only acting as sensors, but also providing testbeds, alongside NV-based quantum sensing to probe fundamental questions about quantum properties of radical pairs and the role of CISS [271, 54, 272].